OpenStudy (anonymous):

a triangle has a height of 15 inches, and an area of 120 inches squared. What is the length?? Please help!! I know that it's late, but this was due today!

jigglypuff314 (jigglypuff314):

area of triangle = (1/2) (base) (height) plug in what you know 120 = (1/2) (base) (15) solve for base

jigglypuff314 (jigglypuff314):

so base = 120 * 2 / 15

OpenStudy (anonymous):

I need the length... what does this mean? Sorry- I am a 4th grader- I got put in stupid advanced math

jigglypuff314 (jigglypuff314):

length is a side, base is the side where the height comes from, which is what your teacher was probably asking for :)

OpenStudy (anonymous):

So it would be 16?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Thank you!!!!!!!

jigglypuff314 (jigglypuff314):

your welcome :)